OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation (x-3)^2/16 - (Y+4)^2/9=1 Vertices: (7, -4), (-1, -4); Foci: (-2, -4), (8, -4) Vertices: (-4, 7), (-4, -1); Foci: (-4, -2), (-4, 8) Vertices: (6, -4), (0, -4); Foci: (0, -4), (6, -4) Vertices: (-4, 6), (-4, 0); Foci: (-4, 0), (-4, 6)

OpenStudy (anonymous):

find the eccentricity first.

OpenStudy (anonymous):

how?

OpenStudy (jdoe0001):

so you have $$ \cfrac{(x-h)^2}{a^2}-\cfrac{(y-k)^2}{b^2}=1\\ \cfrac{(x-3)^2a}{4^2}-\cfrac{(y+4)^2}{3^2}=1\\ $ can you tell which is your "a" element? that is, the one that will have the "traverse axis"

OpenStudy (jdoe0001):

shoot

OpenStudy (jdoe0001):

so you have $$ \cfrac{(x-h)^2}{a^2}-\cfrac{(y-k)^2}{b^2}=1\\ \cfrac{(x-3)^2a}{4^2}-\cfrac{(y+4)^2}{3^2}=1\\ $$ can you tell which is your "a" element? that is, the one that will have the "traverse axis"

OpenStudy (jdoe0001):

what the heck I have an "a" there? lemme repaste

OpenStudy (jdoe0001):

so you have $$ \cfrac{(x-h)^2}{a^2}-\cfrac{(y-k)^2}{b^2}=1\\ \cfrac{(x-3)^2}{4^2}-\cfrac{(y+4)^2}{3^2}=1\\ $$ can you tell which is your "a" element? that is, the one that will have the "traverse axis"

OpenStudy (jdoe0001):

anyhow, the hyperbola is opening towards the axis with the POSITIVE sign

OpenStudy (jdoe0001):

can you find the center of the hyperbola?

OpenStudy (jdoe0001):

is pretty much all there GIVEN in the equation you have really

OpenStudy (jdoe0001):

can you see them now? $$ \cfrac{(x-\color{red}{h})^2}{\color{blue}{a}^2}-\cfrac{(y-\color{red}{k})^2}{b^2}=1\\ \cfrac{(x-\color{red}{(3)})^2}{\color{blue}{4}^2}-\cfrac{(y-\color{red}{(-4)})^2}{3^2}=1\\ $$

OpenStudy (anonymous):

is this the answer? Vertices: (-4, 6), (-4, 0); Foci: (-4, 0), (-4, 6)

OpenStudy (jdoe0001):

dunno

OpenStudy (jdoe0001):

to find the vertices, you'd need to move "a" distance from the center of the hyperbola

OpenStudy (jdoe0001):

over the traverse axis

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

can you find the center in your equation? and the "a" element?